OpenStudy (anonymous):

Show how you solve the following two equations: x – 15 = 32 x – 7.2 = 23.1

OpenStudy (anonymous):

@ Squirrels

OpenStudy (squirrels):

(The first equation) First thing you want to do is isolate x. Isolate means get a number/variable etc. by itself. To do that, you want to add 15 by both sides of the equation... x – 15 = 32 + 15 +15 ----------- ?????? So, Ill let you do this part, and ill show you what to do next.

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

um... I need an online calculator please")

OpenStudy (anonymous):

i cant do math without one

OpenStudy (squirrels):

Lol, its really not necessary, -15 + 15 = 0 and 32 + 15 = 47 x – 15 = 32 + 15 +15 ----------- x=47 See what I did there? :P

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

squ?

OpenStudy (squirrels):

sorry, my browser shut down :P anyways for this equation you are trying to find out what number is "x". So what you want to do it isolate x. (isolate means to get the variable by itself on one side of the equation" x – 15 = 32 <-- So to get "x" to look like this "x=?+?" You have to add 15 by both sides. why? because adding the same number to a negative makes it zero, that way you could get "x" by itself. x – 15 = 32 + 15 +15 ---------- Are you getting it now? :P

OpenStudy (anonymous):

yes i am:)

OpenStudy (squirrels):

ok good :D so x – 15 = 32 + 15 +15 ---------- ??????? <------- can you tell me what it is going to look like now? :P

OpenStudy (anonymous):

k:)

OpenStudy (squirrels):

oh wait no theres not.. continue lol

OpenStudy (anonymous):

lol okay

OpenStudy (squirrels):

I cant really explain it better than that, sorry :/

OpenStudy (anonymous):

um...omg

OpenStudy (anonymous):

it fine squ! i got the gist of it:)

OpenStudy (squirrels):

ok i hope so :)

OpenStudy (anonymous):

:) thanks

OpenStudy (squirrels):

welcome :)

OpenStudy (anonymous):

so if i follow your insructions here i can finish the assignment. if not ill ask for helpfrom someone:)

OpenStudy (squirrels):

sure thing :o have a good day

OpenStudy (anonymous):

u too!